OpenStudy (anonymous):

Circle A has center of (5, -2) and a radius of 7, and circle B has a center of (7, -8) and a radius of 35. What steps will help show that circle A is similar to circle B? (5 points) Select one: a. Rotate circle A 90° counterclockwise about the center. b. Translate circle A using the rule (x-2, y+6). c. Reflect circle A over the y-axis. d. Dilate circle A by a scale factor of 5.

OpenStudy (anonymous):

@ganeshie8

mathslover (mathslover):

Well, can you tell me when are the 2 circles said to be similar?

OpenStudy (anonymous):

im not sure thats why i posted it

mathslover (mathslover):

Okay, no problem. See, two circles can be made similar by finding the difference in the centre coordinates .

mathslover (mathslover):

Here, we have coordinates of centre of Circle A = (5,-2) Circle B = (7,-8) So, we will find the diff. in the above coordinates . Can you find the difference?

OpenStudy (anonymous):

no

mathslover (mathslover):

7-5 = ? -8 - (-2) = ?